Systems Staff Meeting Notes 3/5/01

Systems Staff Meeting Notes
March 5, 2001

Present: Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Default Printer Warning - if you take a printer off the network, you may want to be sure to pick a defined printer for it in replacement - Bill took his home printer off and it unintentionally defined the one here at work and printed to it.
  2. Cracks in the building - new cracks in the wall in the vault and elsewhere have been spotted. Call placed to Facilities and they will contact someone to come and take a look at it this week. Suspect that the underground pipe removal may have something to do with it.
Review Minutes of 2/26/01
  1. Re: Mark Patrick (Health Science Center/PAULA) - Mark was on sabbatical but he is back now and is looking into purchasing a SQL server to run PAULA. We probably won't hear from them until they have it up and running, but it seems to be a-go.
  2. Re: Borland Kylix Server Developer edition software - it has not been received yet, they pre-announced its release date - won't be shipped until late March.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. SuSE Linux 7.1 Professional update not received yet.
  2. 30 more Dells should be arriving this week. Where to store them? Martha has allocated 10 of the 30 to Resource Services. That should provide some space in the room 10 vault. Will probably need to use the 2nd vault for storage as well. We will need to have allocated and removed all of these recently ordered machines before we buy any new ones. We may go back to buying Gateway for the next order - if we make another order before the end of the fiscal year.
Support Issues and Liaisons
  1. Gary Cornwell offered the use of their internal HW tracking database to the rest of the library departments. He will try it in HSS and see how it goes. If it goes over, we would need to convert it to SQL, it is currently in ACCESS.
Networking
  1. CORE problems last week, none of them had too big an effect on us.
  2. Wiring abatement still ongoing. Systems is finished. West 1st floor is done.
  3. Digital is working on wiring to the Swamp and room 148. 104 pulls to be done.
  4. USI returned with solvent to remove the glue on the floor of the West classroom. Did a pretty good job, may not have to get new carpeting after all.
  5. Room 273 - conference room on hold: Facilities needs to re-carpet, Digital can't put it the brass plates in the floor until that is done. The carpeting can't be installed until Dale returns to make a decision on replacing the ceiling tiles. Go figure.
  6. HSS room 100 & 148 - Bobbie was thinking of taking it down temporarily tomorrow. The current rack won't hold; she'd like to put in a taller rack. That would make 11 switches with a total of 220 ports. The math is adding up for Bill, he'd like to take a look at it before committing to that plan.
  7. Journalism - no date scheduled for wiring work yet. Waiting on Facilities to patch the closets. Facilities will also need to move the lights in the ceiling before wiring can begin. Should happen within the next two weeks.
  8. MSL - finishing up.
  9. Need to get a count of UPS's for the closets so we can get more ordered.
  10. Wireless - 5th floor still not done, waiting on parts.
  11. SID server on NT8 - Erich is anxious about his set-up.
Software Projects Status
  1. Trouble Tickets - finished. Bill will inform the liaisons of the changes.
  2. OPS Task Tracker modifications - Hilton working on it.
  3. RS PAULA - Cynthia is still working on enhancements.
  4. ILL Borrowing/Lending tracking system - they currently use CLIO for statistics. They are looking to change to ILLIAD on the SQL server. This would eliminate double entry currently necessary in CLIO. Winston is checking on the necessary technical information and requirements. Looks like it will be acquired.
  5. Personnel database - Winston is incorporating some modifications for Melody and will be re-writing the OPS side. Ying - working on Ringling database, it will be ready for review by Erich on Wednesday.


Back to the Systems Homepage

Last updated March 12, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu